Mar 24, 2016 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Decide on how to implement George's extraction of v_eff constants and use in the analysis.
  2. Update BCAL clusterizer to improve agreement between simulated and generated MC (Mark)
  3. Modify covariance matrix for FCAL and BCAL to correspond to the current resolution of the detectors (Matt).
  4. Change F250 simulation code to match firmware and add flexibility for new algorithms (Mike)


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. BCAL Action Items and Working Tasks

Tentative Agenda

  1. Announcements
  2. Action Items:
  3. Run Preparations
    1. FCAL Update
    2. BCAL Update
  4. Reconstruction and Simulation
  5. Calibration
    1. mcsmear 20160323
    2. BCAL - Effective Velocities and Curvature - George
    3. Instabilities in the Integration of BCAL Pulses - Andrei
  6. Any other business


Attendees: Elton, Mark, Wil, Jon, Adesh, Simon (Jlab); George, Christine (Athens); Mike (CMU); Ahmed, Zisis, Tegan, Andrei (UofR); Sean (NW). Mike (CMU)

  1. Announcements
    1. Accelerator has been down. Leak was found in the refrigerator and they are working to get the system back up for the weekend
    2. Decision for an extension is being debated. Note added: The run has been extended till Mar 25
    3. The cryogenic group found that one of the compressors for the solenoid was problematic and needed to change it out for the spare (there are 2 compressors in use and one spare). It was decided to ramp the magnet down to 800 A before the change over. However, when the compressor was changed out, the solenoid ramped down and is currently at zero current.
  2. Action Items:
  3. Run Preparations
    1. FCAL Update
      1. Base replacement
        • Replaced 13 bases (10 did not produce HV, 3 of them produced 50 V below the set point and beginning to fail)
        • The HV has been turned off for now to save the bases
        • We are working with Chris and Nick to develop a base test and repair station at JLab. We will meet Fri afternoon to see how to move that forward.
      2. Calibration
        • Latest gains have been uploaded to ccdb
        • Updated the HV based on the measured gains. Left the inner circle unmodified out of concerns that gain determination was poor
        • Skim files are now produced during the offline launches, and should help cut down the iteration time for gain determination
      3. Timing
        • Adesh is trying to get timing offsets using neighbors in common showers, but there is a computational challenge.
        • Pedestal plugin can be run over skim files.
        • Sean: The plugin will start running as part of standard plugins once included into the list
      4. Shower split-offs (Jane)
        • Working with Jane to understand the split-offs that are seen in the exclusive data.
        • Ahmed worked with her to investigate some of this behavior using the event viewer hdview2. The hope is to identify the problem and fix this in the clusterizer.
    2. BCAL Update
      • We plan to cycle through various BCAL LED configurations during the production runs.
      • Ahmed took data while he was here to allow him to determine which voltage settings are needed to cover all channels.
      • Ahmed has not yet been able to complete the analysis
      • Orlando Soto is here for shifts and will be working on the script to cycle through various LED configurations.
  4. Reconstruction and Simulation
    1. mcsmear 20160323
      • Sean has produced plots that compare data with MC.
      • For a first look there is reasonable agreement, but several areas in both the BCAL and FCAL where discrepancies need to be understood.
      • There was discussion on what could be the source of the discrepancies. Elton noted that the trigger simulation in MC may be one source of the difference.
  5. Calibration
    1. BCAL - Effective Velocities and Curvature - George
      • veff parameter was extracted for both data and MC
      • Fits were also made to a linear (veff) + quadratic (p2) term.
      • Muon data and comics show little dependence of the linear term veff with layer number
      • For hadrons and production data the quadratic fits, the linear term veff dependence is much reduced, and the quadratic term takes up the layer dependence. This is expected if this is due to the shower size as it spreads out.
      • However, it is interesting to note that the p2 parameter gains a phi dependence.
      • MC: At present the MC uses the ccdb parameters to generate the pulses. Mark: This may not be desirable if the parameters include size-dependent effects that are already included in the MC in other ways.
      • Mark: Shower (curvature) corrections may be double-counting for some of these effects, so they need to be understood and a consistent method developed for their use.
      • Zisis: Measurements of veff at TRIUMF for an early prototype (PolyHitec fibers instead of Kuraray) gave a value of 16.2 cm/ns
    2. Clusterizer (Will and Mark)
      • Will has been working on the cluserizer, evaluating the algorithms on single events with hdview2. Using both data and MC
      • Changes to parameters have improved the visual assignment of hits to the clusters.
      • They are now checking MC data for single photons and also pi0s with higher statistics to stress the new clusterizer algorithms.
      • The progress will be summarized soon in a log book entry
    3. Instabilities in the Integration of BCAL Pulses - Andrei
      • Understanding the distortions in 60-sample and 26-sample data
      • Incorrectly subtracted pedestals is one effect and after pulsing is another, which acts in the opposite direction.
      • There is some compensation for the 60-sample data.
      • He expects to be able to correct the 26-sample data with peak information, which will be available with the next firmware upgrade.
    4. Firmware emulation (Mike)
      • Firmware emulation is complete
      • Software now saves both raw and simulated information in raw data objects, so comparison between variations is easy
      • Comparison of mode 8 data with firmware output shows good agreement.
      • Sean: Noted that of the mode 8 data files that have been written, most of them showed poor FDC efficiencies in the REST files.. Only the last one (run 10836) looked ok. The problem with earlier runs may be calibration.
  6. Any other business